OpenStudy (jas.2002):

Which set of side lengths defines a right triangle? A. 7, 9, 13 B. 5, 12, 14 C. 8, 15, 17 D. 9, 12, 14

Mehek (mehek14):

Use the Pythagorean Theorem

OpenStudy (jas.2002):

but how? im not good at it :/

Mehek (mehek14):

The 2 smaller sides squared must equal to the bigger side squared \(3^2+4^2=5^2\) is an example

OpenStudy (jas.2002):

is it A?

Mehek (mehek14):

\(7^2=49\\9^2 = 81\\49+81 = 130\\ \sqrt{130}=11.4\) So no

OpenStudy (jas.2002):

is it D?

Mehek (mehek14):

\(9^2=81\\12^2=144\\18+144=132\\ \sqrt{132}=11.5\) No

OpenStudy (jas.2002):

omg I dont know.. B?

Mehek (mehek14):

25 + 144 = 169 \(\sqrt{169}=13\)

OpenStudy (jas.2002):

so is that a yes ? Idk im really bad at this stuff..

Mehek (mehek14):

Is 5, 12, and 13 an option?

OpenStudy (jas.2002):

there is 5, 12, and 14..

Mehek (mehek14):

So no

OpenStudy (jas.2002):

so its c?

Mehek (mehek14):

Well yeah because all the other options are wrong

OpenStudy (jas.2002):

haha yeah.. I guess thats a way to do it. thank you so much for all the help

Mehek (mehek14):

yw

OpenStudy (jas.2002):

wow I didnt pass my assessment.. :(